OpenStudy (usukidoll):

Suppose that L is a linear differential operator and L(cosx) doesn't equal to 0. and L(sinx) doesn't equal to 0. Prove that the equation Ly=cosx has a solution of the form A cosx + B sinx

OpenStudy (usukidoll):

@hartnn

OpenStudy (usukidoll):

so... hmm is there a complex root involved or...

OpenStudy (usukidoll):

if I just plug in cosx to Ly=cosx I get L(cosx) = cosx for sinx to Ly=sinx L(sinx) = sinx

OpenStudy (usukidoll):

so... L(cosx) = A L(sinx) =B A = cos x B = sin x A+ B cos x + sin x? my guess

OpenStudy (usukidoll):

@wio

OpenStudy (anonymous):

Linear differential operator means it takes a function and turns it into a linear nth order differential expression, right?

OpenStudy (anonymous):

What about induction on n?

hartnn (hartnn):

try converting cos into its complex form, \(\Large \cos x=\dfrac{e^{ix}+e^{-ix}}{2}\)

OpenStudy (usukidoll):

:/ that looks like the hyperbolic meaning remember cosh?>

OpenStudy (usukidoll):

induction on n?

hartnn (hartnn):

\(\Large \cos hx=\dfrac{e^{x}+e^{-x}}{2}\)

OpenStudy (anonymous):

Like if \(L\) is an \(n\)th order linear differential operator?

OpenStudy (usukidoll):

like for example let's have it at Ly would it be the ...oh no it's going to turn nasty

OpenStudy (usukidoll):

do I convert sinx into the sinhx complex form too?

OpenStudy (anonymous):

If you start with \(n=0\) then you have: \[ a_1y=\cos x \]

OpenStudy (anonymous):

Leting \(A=\frac 1{a_1}\) and \(B=0\) this is true.

OpenStudy (anonymous):

Then try to prove for \(n\) assuming that \(n-1\) and \(n-2\) etc would work.

OpenStudy (anonymous):

Also, hartnn wasn't using coshx and sinhx he's using the euler definition cosx + isinx = e^ix to rewrite cosx and sinx in terms of exponentials.

OpenStudy (anonymous):

So you have \[ a_ny^{(n)}+L_2(y) = \cos x \]Where \(L_2(y)\) is some \(n-1\)th order linear differential operator, and we can assume by induction it has a solution in the form \(A\cos x +B\sin x\)

OpenStudy (anonymous):

@UsukiDoll Are you understanding my induction here?

OpenStudy (usukidoll):

oh yeah

OpenStudy (usukidoll):

I haven't gotten to Euler's version of sinx and cosx, but I still remember euler's method

OpenStudy (usukidoll):

sorry was doing my english assignment...reading my model writer's work

OpenStudy (usukidoll):

so I should do the same thing for sin x... an y^n+L1y = sin x?

OpenStudy (anonymous):

You only have to do it for \(Ly=\cos x\) though... right?

OpenStudy (usukidoll):

oh yeah right... so no need for sin x

OpenStudy (usukidoll):

so for n-1 An-1y^(n-1)+L2(y)=cos x